# Exploring Test-Time Adaptation for Object Detection in Continually Changing Environments

### v4

@@ 3.2.1. Overview
Figure 2 presents a concise overview of AMROD, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Adaptive Randomized Restoration (ARR). Inspired by UDA object detection methods (Li et al., 2022a; Liu et al., 2021), we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student for gradient descent is defined as:
@@
where denotes the contrastive learning (CL) loss, to be elaborated later. The represents the Kullback-Leibler (KL) Divergence (Kullback and Leibler, 1951) loss, used to quantify the distinction between two probability distributions. The parameters and are corresponding weights. KL loss is defined as:
@@
3.2.2. Object-Level Contrastive Learning
SimCLR is a widely used CL approach for self-supervised visual representation learning (Chen et al., 2020). It learns high-quality feature representations from differently augmented views of the same image, which requires large batch sizes to ensure sufficient positive and negative pairs for classification tasks. Consequently, SimCLR is less suitable for object detection, where images contain multiple instances and large batch sizes demand significant computational resources.
Motivated by SimCLR (Chen et al., 2020), we present an OCL module to extract teacher and student features for CL based on proposals from the RPN. This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detector updates. Specifically, given a weakly augmented image at time step t, the teacher produces ROI proposals via region proposal network. OCL then apply RoIAlign (He et al., 2017) to extract corresponding teacher and student object-level features and based on the feature map from the backbone, respectively. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. Then CL loss is applied to these features and derived from weak and strong enhanced images respectively by minimizing:
where is the temperature, indicates the number of the features, and and denote the features of two different augmentations of the same object, serving as the positive pair. This strategy encourages the model to learn fine-grained and localized feature representations on the target domain, without relying on accurate pseudo-labels.
3.2.3. Adaptive Monitoring
@@
3.2.4. Adaptive Randomized Restoration
@@
Existing methods address catastrophic forgetting through stochastic reset (Wang et al., 2022) or pure data-driven restoration (Brahma and Rai, 2023). Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the most unimportant parameters, which may retain noise parameters, thereby leading to error accumulation. Unlike PETAL (Brahma and Rai, 2023), which was designed for classification tasks and resets the least active parameters across all layers, our approach refines this for object detection by identifying inactive parameters within each layer individually. This layer-specific technique accounts for the varying parameter scales at different depths of the model, and by further incorporating randomness, we propose ARR to reset irrelevant parameters with higher possibilities to enhance stability and robustness in restoration. Following (Brahma and Rai, 2023), we adopt the diagonal approximation of Fisher Information Matrix (FIM) (Kirkpatrick et al., 2017) based on gradient, as a metric of parameter importance. The FIM approximates parameter importance by considering the sensitivity of the loss function to parameter changes (Kirkpatrick et al., 2017). Parameters with higher values in the FIM are those for which small changes result in significant increases in loss, indicating their importance for the domain at hand.
Specifically, given a layer with weight updated through gradient descent to based on gradient matrix at time step , ARR generates a random matrix with the same shape as the weights matrix, where each element follows a uniform distribution between 0 and 1. Then the element-wise multiplication of the squared gradient matrix and the random matrix is applied to approximate the FIM. Furthermore, ARR updates the parameters based on the FIM values, formulated as:
@@
where is the element-wise multiplication, indicates the mask matrix, represents element-wise less than operation, and is the threshold value acquired by the -quantile of : . Consequently, the elements in are set to 1, when the corresponding FIM value is less than , indicating the corresponding parameters should be reset to the source weight . This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness to mitigate forgetting.
4. EXPERIMENTS
@@
CoTTA | 7.8 | 9.0 | 8.9 | 1.8 | 0.3 | 7.1 | 38.4 | 31.1 | 8.6 | 49.6 | 16.2 | 13.1 | 16.0 | +0.9 | 6k |
SVDP | 7.7 | 10.1 | 9.7 | 2.3 | 0.7 | 13.0 | 42.4 | 45.2 | 15.4 | 47.2 | 21.2 | 14.8 | 19.1 | +4.0 | 6k |
IRG | 8.0 | 11.0 | 9.3 | 3.4 | 1.2 | 13.0 | 37.9 | 41.3 | 15.9 | 38.9 | 16.9 | 13.4 | 17.5 | +2.4 | 6k |
MemCLR | 8.5 | 10.4 | 10.6 | 2.7 | 1.1 | 12.2 | 41.4 | 41.6 | 16.4 | 43.1 | 15.4 | 12.7 | 18.0 | +2.9 | 6k |
Ours-unstop | 8.6 | 12.1 | 11.7 | 3.6 | 1.5 | 16.7 | 44.7 | 48.1 | 16.7 | 47.4 | 22.5 | 13.9 | 20.6 | +5.5 | 6k |
@@
CoTTA | 38.2 | 10.6 | 0.4 | 33.8 | 9.3 | 40.8 | 10.4 | 0.6 | 36.5 | 9.2 | 40.8 | 10.4 | 0.5 | 36.3 | 9.6 | 19.1 | +2.7 | 25k |
SVDP | 36.8 | 8.8 | 0.6 | 43.5 | 11.2 | 45.0 | 15.4 | 3.8 | 47.8 | 19.9 | 41.5 | 16.2 | 4.6 | 44.6 | 20.2 | 25.3 | +8.9 | 25k |
IRG | 37.9 | 9.0 | 0.7 | 44.3 | 12.2 | 45.6 | 17.7 | 6.0 | 46.7 | 21.9 | 37.0 | 16.7 | 7.1 | 38.5 | 20.5 | 25.6 | +9.2 | 25k |
MemCLR | 37.7 | 8.9 | 0.8 | 45.5 | 13.5 | 45.0 | 18.1 | 5.1 | 46.8 | 22.4 | 36.8 | 18.5 | 7.5 | 38.4 | 21.8 | 26.0 | +9.6 | 25k |
Ours-unstop | 39.0 | 10.4 | 0.8 | 48.0 | 13.8 | 49.4 | 18.6 | 7.7 | 51.7 | 25.0 | 45.7 | 20.0 | 11.8 | 46.4 | 25.8 | 29.0 | +12.6 | 25k |
Ours | 39.2 | 9.9 | 0.8 | 46.7 | 13.8 | 49.4 | 18.5 | 8.0 | 52.9 | 24.7 | 46.2 | 19.5 | 11.3 | 47.3 | 29.1 | 29.2 | +12.8 | 20.1k |
In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection. These tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA (Wang et al., 2022), the short-term CTTA task entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA task involves continually adapting the model toward a group of target domains cyclically.
4.1. Datasets
Cityscapes. The Cityscapes dataset (Cordts et al., 2016) is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight object types.
We utilize the model pre-trained on the training set of Cityscapes as the source model, and the source data is discarded during adaptation.
Cityscapes-C. The Cityscapes-C (Hendrycks and Dietterich, 2019) is initially designed to assess robustness against various corruptions, introducing 15 types of corruption with 5 levels of severity. We create the dataset by applying these corruptions at the maximum severity level 5 to the validation set of clean cityscapes, treating each corruption type as an individual target domain comprising 500 images.
Our short-term CTTA task selectively focuses on the latter 12 corruptions.
For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain group following (Gan et al., 2023b), namely Fog, Motion, Snow, Brightness, and Defocus.
@@ 4.1. Datasets
SHIFT. The SHIFT (Sun et al., 2022) is a synthetic dataset for autonomous driving, featuring real-world environmental changes. SHIFT can be categorized as clear, cloudy, overcast, rainy, and foggy, where each condition contains images taken at various times ranging from daytime to night.
@@
ACDC. The ACDC (Sakaridis et al., 2021) shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following (Wang et al., 2022), we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition. Similarly, the source model is continually adapted to the target domain group for 10 cycles.
4.2. Implementation Details
We adopt the Faster R-CNN with ResNet50 (He et al., 2016) pre-trained on ImageNet (Krizhevsky et al., 2012) as the backbone. Aligning with (VS et al., 2023b; Wang et al., 2022), we maintain a batch size of 1 to emulate a real-world application scenario where the detection model adapts toward a continuous influx of images. The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform (Wu et al., 2019). The metric of mAP at an IoU threshold of 0.5 (mAP@0.5) is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More hyper-parameters and implementation details are provided in the supplementary material.
4.3. Baselines and Compared Approaches
We compare AMROD with six baselines across various settings, including Source (Ren et al., 2015), IRG (VS et al., 2023a), Memclr (VS et al., 2023b), CoTTA (Wang et al., 2022), SVDP (Yang et al., 2024), and AMROD-upstop. Specifically, ”Source” represents the source model without adaptation. Memclr and IRG are SFDA object detection methods, also referred to as TTA methods (Wang et al., 2022). MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL. CoTTA utilizes a fixed threshold for pseudo-labeling and random neuron recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction. “AMROD-unstop” represents AMROD without dynamic skipping.
4.4. Results and Analysis
4.4.1. Synthetic Continual Distribution Shift
We first evaluate the effectiveness of AMROD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks. As depicted in Table 1, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions. Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches. Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to AMROD-unstop. AMROD ensures a more reliable and efficient adaptation to target domains characterized by intense and continual changes in the short term.
@@ 4.4.1. Synthetic Continual Distribution Shift
CoTTA | 53.7 | 19.7 | 38.0 | 42.4 | 53.1 | 19.7 | 37.7 | 42.9 | 51.3 | 19.0 | 36.5 | 41.8 | 50.9 | 19.1 | 36.0 | 42.4 | 37.8 | +1.8 | 16k |
SVDP | 52.8 | 20.0 | 35.6 | 42.0 | 54.6 | 23.5 | 38.7 | 43.8 | 52.8 | 24.0 | 38.6 | 43.9 | 51.8 | 23.6 | 38.2 | 43.0 | 39.5 | +3.5 | 16k |
IRG | 52.7 | 20.6 | 36.0 | 42.9 | 53.6 | 23.2 | 38.1 | 45.6 | 51.0 | 23.1 | 37.1 | 42.8 | 49.7 | 22.5 | 36.5 | 40.7 | 38.9 | +2.9 | 16k |
MemCLR | 52.9 | 21.2 | 35.2 | 42.8 | 52.8 | 23.2 | 38.2 | 44.4 | 51.4 | 23.2 | 37.1 | 42.9 | 49.9 | 22.9 | 36.6 | 41.0 | 38.7 | +2.7 | 16k |
Ours-unstop | 52.9 | 20.4 | 34.5 | 42.9 | 54.5 | 23.8 | 39.6 | 45.8 | 53.5 | 25.1 | 39.0 | 45.3 | 52.5 | 24.3 | 38.8 | 44.7 | 40.3 | +4.3 | 16k |
@@
As presented in Table 8, the long-term task outcomes reveal the source model’s poor performance, with an average mAP of 16.4. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP. Particularly, AMROD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the ARR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
@@
4.4.2. Real-World Continual Distribution Shift
We also evaluate our method on the real-world continual distribution shift dataset. The results of the SHIFT short-term CTTA task are shown in Table 4. While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling, stochastic restoration, and prompts learning achieves a 0.5 mAP improvement. Moreover, AMROD which introduces AM and ARR modules shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
The experimental results of the ACDC long-term CTTA task are summarized in Table 3. CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, AMROD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines. While AMROD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
Mean-Teacher | OCL | AM | ARR | Mean | Gain | |
1 | 25.9 | / | ||||
2 | 26.7 | +0.8 | ||||
3 | 27.3 | +1.4 | ||||
4 | 27.4 | +1.5 | ||||
5 | 28.6 | +2.7 | ||||
6 | 29.2 | +3.3 |
4.4.3. Qualitative Results

### v5

@@ III-B1 Overview
Figure 2 presents a concise overview of AMROD with details described in Algorithm 1, consisting of Object-level Contrastive Learning (OCL), Adaptive Monitoring (AM), and Adaptive Randomized Restoration (ARR). Inspired by UDA object detection methods [37, 79], we adopt the Weak-Strong augmentation to enable the teacher model to generate reliable pseudo-labels without being affected by heavy augmentation. Specifically, the teacher receives input with weak augmentations while we input student networks with strong enhanced images. The overall loss of the student for gradient descent is defined as:
@@
where denotes the contrastive learning (CL) loss, to be elaborated later. The represents the Kullback-Leibler (KL) Divergence [80] loss, used to quantify the distinction between two probability distributions. The parameters and are corresponding weights. KL loss is defined as:
@@
III-B2 Object-Level Contrastive Learning
We first briefly introduce SimCLR, a widely used CL approach for self-supervised visual representation learning [81]. SimCLR aims to learn high-quality feature representation across differently augmented views of the same image. For the image in a batch, the CL loss in SimCLR is formulated as follows:
where is a temperature hyper-parameter. and denote the features of two different augmentations of the same images, serving as the positive pair. At the same time, represents one of the features of a different augmented sample in a batch size of . Here, the indicates the similarity function, i.e. cosine similarity. Notably, the SimCLR is initially designed for classification tasks, assuming each image pertains to a single category, and requires large batch sizes to ensure sufficient positive and negative pairs for feature representation learning. Consequently, the original SimCLR is not well-suited for object detection, where images typically contain multiple instances, thus requiring significant computational resources to accommodate large batch sizes.
Motivated by SimCLR [81], we present an OCL module to extract teacher and student features for CL based on proposals from the RPN. This strategy provides multiple cropped views around the object instance, eliminating the need for large batch sizes and ensuring computational efficiency for online detector updates. Specifically, given a weakly augmented image at time step t, the teacher produces ROI proposals via region proposal network. OCL then apply RoIAlign [82] to extract corresponding teacher and student object-level features and based on the feature map from the backbone, respectively. The features associated with the same proposal are considered positive pairs, otherwise negative pairs. Then CL loss is applied to these features and images respectively by minimizing:
where is the temperature, indicates the number of the features, and and denote the features of two different augmentations of the same object, serving as the positive pair. This strategy encourages the model to learn fine-grained and localized feature representations on the target domain, without relying on accurate pseudo-labels. Moreover, the OCL is well integrated into the mean-teacher self-training paradigm as a drop-in enhancement for feature adaptation.
III-B3 Adaptive Monitoring
@@
III-B4 Adaptive Randomized Restoration
@@
Long-term adaptation to dynamic scenarios can lead to catastrophic forgetting, where self-training may reinforce erroneous predictions. Even worse, the model might fail to recover when transitioning to a new domain. Therefore, addressing the forgetting issue is especially crucial for CTTA. Existing methods address catastrophic forgetting through stochastic reset [24] or pure data-driven restoration [49]. Nevertheless, the former may reset valuable parameters, potentially erasing essential knowledge relevant to the current domain. On the other hand, the latter restores the least important parameters, which may retain noise parameters, thereby leading to error accumulation. Moreover, PETAL [49], designed for classification tasks, resets the least active parameters globally across the entire network. This global strategy can overlook the non-uniform scales of weights in different layers, particularly in deeper neural networks. Our approach refines this for object detection by identifying inactive parameters within each layer individually. This layer-specific technique accounts for varying parameter scales at different model depths. Building upon this, our proposed ARR mechanism further incorporates randomness to reset irrelevant parameters with a higher probability, thereby enhancing stability and robustness during restoration.
Let denote the distribution over model prediction , parameterized by , given an input . The impact of a small perturbation to these parameters can be quantified by the KL divergence between the original and perturbed distributions. [83, 84] shows that as , the following second-order approximation holds:
where is the Fisher information matrix (FIM) [31, 32], defined as:
This approximation demonstrates that the FIM links parameter perturbations to the resultant changes in the model’s output distribution, indicating their importance. Consequently, we leverage FIM to guide the parameter restoration process. However, the dimensionality of FIM makes it intractable to compute in practice. Therefore, consistent with prior work [76], we adopt the diagonal approximation of FIM, represented as a vector in . Moreover, in practice, machine learning models are typically trained on a finite set of N training samples , rather than having direct access to the true data distribution . In such case, the diagonal FIM can be empirically approximated:
where is the empirical diagonal FIM. Each component of corresponds to an individual parameter, and a larger value for such a component signifies greater influence of the corresponding parameter on the model’s predictions.
For CTTA, given a batch of pairs of unlabeled test inputs and corresponding teacher’s pseudo-label at time step , the FIM for the student model is defined as:
where denotes the student model parameters after gradient descent from at time step . Since the FIM is derived from pseudo-labels, which can be noisy in the target domain, the FIM itself might assign high importance to parameters associated with these erroneous predictions. Therefore, given a specific layer with parameters (still denoted as for simplicity), ARR integrates a stochastic competent by employing a random matrix with the same shape as the parameters, where each element follows a uniform distribution between 0 and 1. The reset scores are then obtained by an element-wise multiplication of the FIM and the random matrix. For each layer, the parameters of the student model are updated through:
@@
where is the element-wise multiplication, indicates the mask matrix, represents element-wise less than operation, and is the threshold value acquired by the -quantile of : . Consequently, the elements in are set to 1, when the corresponding score value is less than , indicating the associated parameters should be reset to the source parameters . This strategy enables the model to retain essential knowledge while introducing randomness to enhance robustness to mitigate forgetting.
IV EXPERIMENTAL SETUPS
@@
Tent | 6.8 | 7.8 | 7.7 | 1.3 | 0.2 | 6.1 | 33.1 | 28.0 | 2.2 | 51.1 | 14.8 | 11.0 | 14.2 | -0.9 | 6.0k |
CoTTA | 7.8 | 9.0 | 8.9 | 1.8 | 0.3 | 7.1 | 38.4 | 31.1 | 8.6 | 49.6 | 16.2 | 13.1 | 16.0 | +0.9 | 6.0k |
SVDP | 7.7 | 10.1 | 9.7 | 2.3 | 0.7 | 13.0 | 42.4 | 45.2 | 15.4 | 47.2 | 21.2 | 14.8 | 19.1 | +4.0 | 6.0k |
IRG | 8.0 | 11.0 | 9.3 | 3.4 | 1.2 | 13.0 | 37.9 | 41.3 | 15.9 | 38.9 | 16.9 | 13.4 | 17.5 | +2.4 | 6.0k |
MemCLR | 8.5 | 10.4 | 10.6 | 2.7 | 1.1 | 12.2 | 41.4 | 41.6 | 16.4 | 43.1 | 15.4 | 12.7 | 18.0 | +2.9 | 6.0k |
Ours-unstop | 8.6 | 12.1 | 11.7 | 3.6 | 1.5 | 16.7 | 44.7 | 48.1 | 16.7 | 47.4 | 22.5 | 13.9 | 20.6 | +5.5 | 6.0k |
@@
Tent | 35.8 | 8.1 | 0.2 | 29.2 | 6.2 | 30.7 | 10.2 | 0.8 | 27.9 | 12.4 | 20.4 | 5.0 | 0.1 | 11.5 | 2.4 | 12.0 | -4.4 | 25.0k |
CoTTA | 38.2 | 10.6 | 0.4 | 33.8 | 9.3 | 40.8 | 10.4 | 0.6 | 36.5 | 9.2 | 40.8 | 10.4 | 0.5 | 36.3 | 9.6 | 19.1 | +2.7 | 25.0k |
SVDP | 36.8 | 8.8 | 0.6 | 43.5 | 11.2 | 45.0 | 15.4 | 3.8 | 47.8 | 19.9 | 41.5 | 16.2 | 4.6 | 44.6 | 20.2 | 25.3 | +8.9 | 25.0k |
IRG | 37.9 | 9.0 | 0.7 | 44.3 | 12.2 | 45.6 | 17.7 | 6.0 | 46.7 | 21.9 | 37.0 | 16.7 | 7.1 | 38.5 | 20.5 | 25.6 | +9.2 | 25.0k |
MemCLR | 37.7 | 8.9 | 0.8 | 45.5 | 13.5 | 45.0 | 18.1 | 5.1 | 46.8 | 22.4 | 36.8 | 18.5 | 7.5 | 38.4 | 21.8 | 26.0 | +9.6 | 25.0k |
Ours-unstop | 39.0 | 10.4 | 0.8 | 48.0 | 13.8 | 49.4 | 18.6 | 7.7 | 51.7 | 25.0 | 45.7 | 20.0 | 11.8 | 46.4 | 25.8 | 29.0 | +12.6 | 25.0k |
Ours | 39.2 | 9.9 | 0.8 | 46.7 | 12.7 | 49.4 | 18.5 | 8.0 | 52.9 | 24.7 | 46.2 | 19.5 | 11.3 | 47.3 | 29.1 | 29.2 | +12.8 | 20.1k |
In this study, we rigorously evaluate our methodology across four benchmark tasks in CTTA object detection. These tasks encompass continual adaptation to synthetic and real-world distribution shifts, evaluated over short and long-term periods. Inspired by the foundational work CoTTA [24], the short-term CTTA tasks entails the sequential adaptation to various target domains once. In contrast, the long-term CTTA tasks involves continually adapting the model toward a group of target domains cyclically.
IV-A Datasets
Cityscapes. The Cityscapes [33] is collected for urban scene understanding, encompassing 2,975 training images and 500 validation images with eight categories, i.e. person, rider, car, truck, bus,
train, motorcycle, and bicycle.
We utilize the model pre-trained on the training set of Cityscapes as the source model, and source data is discarded during adaptation.
Cityscapes-C. The Cityscapes-C [11] is initially designed to assess robustness against various corruptions, introducing 15 types of corruption with 5 severity levels. We create the dataset by applying these corruptions at the maximum severity level to the validation set of clean cityscapes, treating each corruption as an individual target domain comprising 500 images.
Our short-term CTTA tasks selectively focuses on the latter 12 corruptions, including Defocus Blur, Frosted Glass Blur, Motion Blur, Zoom Blur, Snow, Frost, Fog, Brightness, Contrast, Elastic, Pixelate, and JPEG.
For the long-term CTTA task, we prioritize the five corruptions related to autonomous driving scenarios as the target domain group following [66], namely Fog, Motion, Snow, Brightness, and Defocus.
@@ IV-A Datasets
SHIFT. The SHIFT [34] is a synthetic dataset for autonomous driving, featuring real-world environmental changes. SHIFT can be categorized as clear, cloudy, overcast, rainy, and foggy, where each condition contains images taken at various times ranging from daytime to night.
@@
ACDC. The ACDC [35] shares the same class types as Cityscapes and is collected in four different adverse visual conditions, including Fog, Night, Rain, and Snow. Following [24], we use these four conditions as the target domain group for the long-term CTTA task, with 400 unlabeled images per condition. Similarly, the source model is continually adapted to the target domain group for 10 cycles.
IV-B Implementation Details
We adopt the Faster R-CNN with ResNet50 [1] pre-trained on ImageNet [85] as the backbone. Aligning with [16, 24], we maintain a batch size of 1 to emulate a real-world application scenario where the detection model adapts toward a continuous influx of images. The source models are trained using an SGD optimizer with a learning rate of 0.001 and a momentum of 0.9. Algorithms are implemented leveraging the Detectron2 platform [86]. The metric of mAP at an IoU threshold of 0.5 (mAP@0.5) is employed for evaluation. Each experiment is conducted on 1 NVIDIA A800 GPU. More hyper-parameters and implementation details are provided in the supplementary material.
IV-C Baselines and Compared Approaches
We compare AMROD with seven source-free baselines across various settings, including Source [3], Tent [23], IRG [43], MemCLR [16], CoTTA [24], SVDP [87], and AMROD-upstop. Specifically, “Source” represents the source model without adaptation. Tent [23] updates the affine parameters through entropy minimization in TTA. Furthermore, Memclr and IRG are SFDA object detection methods, also referred to as TTA methods [24]. MemCLR integrates cross-attention with CL, while IRG incorporates instance relation graph and CL. CoTTA utilizes a fixed threshold for pseudo-labeling and random neuron recovery to tackle CTTA, while SVDP explores sparse visual prompts for CTTA dense prediction. “AMROD-unstop” represents AMROD without dynamic skipping.
@@ IV-C Baselines and Compared Approaches
Tent | 52.4 | 18.6 | 33.4 | 38.9 | 51.7 | 17.4 | 31.4 | 36.0 | 45.8 | 14.6 | 28.1 | 28.5 | 35.0 | 9.5 | 21.5 | 18.3 | 30.5 | -5.5 | 16.0k |
CoTTA | 53.7 | 19.7 | 38.0 | 42.4 | 53.1 | 19.7 | 37.7 | 42.9 | 51.3 | 19.0 | 36.5 | 41.8 | 50.9 | 19.1 | 36.0 | 42.4 | 37.8 | +1.8 | 16.0k |
SVDP | 52.8 | 20.0 | 35.6 | 42.0 | 54.6 | 23.5 | 38.7 | 43.8 | 52.8 | 24.0 | 38.6 | 43.9 | 51.8 | 23.6 | 38.2 | 43.0 | 39.5 | +3.5 | 16.0k |
IRG | 52.7 | 20.6 | 36.0 | 42.9 | 53.6 | 23.2 | 38.1 | 45.6 | 51.0 | 23.1 | 37.1 | 42.8 | 49.7 | 22.5 | 36.5 | 40.7 | 38.9 | +2.9 | 16.0k |
MemCLR | 52.9 | 21.2 | 35.2 | 42.8 | 52.8 | 23.2 | 38.2 | 44.4 | 51.4 | 23.2 | 37.1 | 42.9 | 49.9 | 22.9 | 36.6 | 41.0 | 38.7 | +2.7 | 16.0k |
Ours-unstop | 52.9 | 20.4 | 34.5 | 42.9 | 54.5 | 23.8 | 39.6 | 45.8 | 53.5 | 25.1 | 39.0 | 45.3 | 52.5 | 24.3 | 38.8 | 44.7 | 40.3 | +4.3 | 16.0k |
@@
V Results and Analysis
V-A Synthetic Continual Distribution Shift
V-A1 Short-term CTTA Tasks Results
We first evaluate the effectiveness of AMROD on the short-term Cityscapes-to-Cityscapes-C adaptation tasks. As depicted in Table II, Tent [23] undergoes a slight decline in performance, dropping from 15.1 to 14.2 mAP relative to the source model. This downturn may be attributed to its dependency on a large batch size to update the parameters of the batchnorm layer, making it suboptimal for online adaptation. In contrast, CoTTA, IRG, and Memclr exhibit enhancements in performance, achieving 16.0, 17.5, and 18.0 mAP respectively. Although the SFDA methods like IRG and MemCLR assume a static target domain, employing a large momentum update rate for the teacher enables a relatively stable adaptation in the short term. Consequently, CL leads to a better performance than CoTTA which solely employs weight-averaged predictions. Moreover, SVDP, which utilizes visual prompts, achieves a sub-optimal performance of 19.1 mAP. Remarkably, our proposed method improves the performance to 20.8 mAP, consistently outperforming the above approaches. Additionally, the dynamic skipping strategy in AM reduces adaptation iterations by 20% while achieving a 0.2 mAP improvement compared to AMROD-unstop. AMROD ensures a more reliable and efficient adaptation to target domains characterized by intense changes in the short term.
V-A2 Long-term CTTA Tasks Results
As presented in Table III, the long-term tasks outcomes reveal the source model’s poor performance, with an average mAP of 16.4. Despite the improvement of Memclr and IRG, their performance also begins to decline in the later rounds, failing to maintain stability over long-term adaptation. We believe this is due to the aforementioned methods not accounting for continual distribution shifts, resulting in error accumulation and catastrophic forgetting. Furthermore, CoTTA utilizes a stochastic restoration mechanism to mitigate forgetting, but its randomness may result in losing crucial information, thus limiting its performance at 19.1 mAP. In contrast, SVDP which employs sparse visual prompts raises the performance to 25.3 mAP. Particularly, AMROD yields a remarkable 12.8 mAP enhancement over the Source and surpasses all comparative baselines using 80% adaptation iterations, which employs the ARR mechanism to conserve valuable knowledge while eliminating noise from prior domains. These findings empirically validate the effectiveness of our proposed method in ensuring stable and efficient adaptation amidst synthetic continual distribution shifts, across both short-term and long-term scenarios.
V-B Real-World Continual Distribution Shift
V-B1 Short-term CTTA Tasks Results
We also evaluate our method on the real-world continual distribution shift dataset. The results of the SHIFT short-term CTTA tasks are shown in Table V. While other baseline methods suffer from performance decline, SVDP which employs a fixed threshold for pseudo-labeling, stochastic restoration, and prompts learning achieves a 0.5 mAP improvement. Moreover, AMROD which introduces the AM module shows a superior performance of 43.9 mAP with 16% improvement in efficiency.
V-B2 Long-term CTTA Tasks Results
The experimental results of the ACDC long-term CTTA tasks are summarized in Table IV. CoTTA, IRG, and MemCLR potentially suffer from error accumulation and catastrophic forgetting, manifesting in a rapid decline in performance during the later stages. Similarly, SVDP achieves a sub-optimal performance of 39.5 mAP. Notably, AMROD achieves a performance of 40.4 mAP, yielding an absolute improvement of 0.9 mAP over baselines. While AMROD may not display a significant advantage over other methods in the initial round, its performance continually improves and remains stable throughout the long-term adaptation. These results underscore the capability of our method to foster robust adaptation toward the real-world continual distribution shift in both the short and long term.
V-C Qualitative Results